TIME FOR A LUMP SUM TO DOUBLE If you deposit money today in an account that pays 10.5% annual interest, how long will it take to double your money? Round your answer to two decimal places.
Answer #1

current value or PV=1

future value or FV=2 as double amount

FV=PV*(1+r)^n

2=1*(1+10.5%)^n

1.105^n=2

take natural log on both sides

n=ln(2)/ln(1.105)

=6.94 years

the above is answer..